OpenStudy (anonymous):

Factor the polynomial expression. x^2-6x-7

OpenStudy (whpalmer4):

Can you think of two numbers which give you -6 if you add, and -7 if you multiply?

OpenStudy (whpalmer4):

Hint: one is positive, the other is negative

OpenStudy (anonymous):

3 and 3?

OpenStudy (whpalmer4):

3 * 3 = -7?!?

OpenStudy (whpalmer4):

@JonyS the idea is to teach HOW to do it, not just supply answers. See the Code of Conduct link at the bottom of the window

OpenStudy (anonymous):

oh should I delete my answer then?

OpenStudy (whpalmer4):

As your penance, please explain how you came up with the factoring...

OpenStudy (anonymous):

@whpalmer4 I think he hasn't seen the answer I'll delete it xD

OpenStudy (anonymous):

If you are struggling with this you can watch this video: http://www.youtube.com/watch?v=eF6zYNzlZKQ

OpenStudy (whpalmer4):

Nah, leave it up there, I'll just add a bit more about why we do that... (X+a)(x+b) = x*x +b*x + a*x + a*b = x^2 + (a+b) + ab Here we have a+b =-6 and a*b=-7 -7*1 =-7 -7+1=-6 So our factoring is (x-7)(x+1)